# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ C.2 Additional Simulation for Transfer Between Different Graphons
$2\rightarrow 1$ 1.9 $\pm$ 0.2 1.8 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$2\rightarrow 3$ 1.6 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$2\rightarrow 4$ 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$2\rightarrow 5$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$2\rightarrow 6$ 1.0 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$2\rightarrow 7$ 1.6 $\pm$ 0.2 1.6 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
$2\rightarrow 8$ 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$2\rightarrow 9$ 1.8 $\pm$ 0.5 1.6 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$\rightarrow 10 2.2$ $\pm$ 0.3 2.2 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$3\rightarrow 2$ 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$3\rightarrow 4$ 1.4 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$3\rightarrow 5$ 1.4 $\pm$ 0.2 1.4 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$3\rightarrow 6$ 0.9 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$3\rightarrow 7$ 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
$3\rightarrow 8$ 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$3\rightarrow 9$ 1.9 $\pm$ 0.3 1.8 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$3\rightarrow 10$ 2.2 $\pm$ 0.3 2.1 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$4\rightarrow 1$ 1.7 $\pm$ 0.2 1.7 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$4\rightarrow 2$ 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$4\rightarrow 3$ 1.4 $\pm$ 0.3 1.3 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$4\rightarrow 5$ 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
$4\rightarrow 6$ 1.0 $\pm$ 0.2 1.1 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$4\rightarrow 7$ 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
$4\rightarrow 8$ 1.4 $\pm$ 0.2 1.4 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$4\rightarrow 9$ 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
$4\rightarrow 10$ 2.2 $\pm$ 0.3 2.1 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
$5\rightarrow 1$ 2.0 $\pm$ 0.3 2.1 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.6 $\pm$ 0.2 2.1 $\pm$ 0.3
$5\rightarrow 2$ 1.7 $\pm$ 0.3 1.6 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$5\rightarrow 3$ 1.7 $\pm$ 0.3 1.7 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$5\rightarrow 4$ 1.4 $\pm$ 0.1 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$5\rightarrow 6$ 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
$5\rightarrow 7$ 1.8 $\pm$ 0.3 1.8 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.2 2.4 $\pm$ 0.6
$5\rightarrow 8$ 1.5 $\pm$ 0.2 1.7 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
$5\rightarrow 9$ 2.1 $\pm$ 0.3 2.0 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.8 $\pm$ 0.3 2.1 $\pm$ 0.1
$6\rightarrow 1$ 2.0 $\pm$ 0.3 2.0 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
$6\rightarrow 2$ 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
$6\rightarrow 3$ 1.5 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
$6\rightarrow 4$ 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
$6\rightarrow 5$ 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0

### v2

@@ E.2 Additional Simulation for Transfer Between Different Graphons
2 $\rightarrow$ 1 1.9 $\pm$ 0.2 1.8 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
2 $\rightarrow$ 3 1.6 $\pm$ 0.3 1.6 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
2 $\rightarrow$ 4 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
2 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
2 $\rightarrow$ 6 1.0 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
2 $\rightarrow$ 7 1.6 $\pm$ 0.2 1.6 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
2 $\rightarrow$ 8 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
2 $\rightarrow$ 9 1.8 $\pm$ 0.5 1.6 $\pm$ 0.4 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
2 $\rightarrow$ 10 2.2 $\pm$ 0.3 2.2 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
3 $\rightarrow$ 1 1.6 $\pm$ 0.5 1.7 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
3 $\rightarrow$ 2 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
3 $\rightarrow$ 4 1.4 $\pm$ 0.2 1.3 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
3 $\rightarrow$ 5 1.4 $\pm$ 0.2 1.4 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
3 $\rightarrow$ 6 0.9 $\pm$ 0.2 1.0 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
3 $\rightarrow$ 7 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
3 $\rightarrow$ 8 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
3 $\rightarrow$ 9 1.9 $\pm$ 0.3 1.8 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
3 $\rightarrow$ 10 2.2 $\pm$ 0.3 2.1 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
4 $\rightarrow$ 1 1.7 $\pm$ 0.2 1.7 $\pm$ 0.2 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
4 $\rightarrow$ 2 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
4 $\rightarrow$ 3 1.4 $\pm$ 0.3 1.3 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
4 $\rightarrow$ 5 1.3 $\pm$ 0.2 1.3 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0
4 $\rightarrow$ 6 1.0 $\pm$ 0.2 1.1 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
4 $\rightarrow$ 7 1.5 $\pm$ 0.2 1.5 $\pm$ 0.2 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.3 2.4 $\pm$ 0.6
4 $\rightarrow$ 8 1.4 $\pm$ 0.2 1.4 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
4 $\rightarrow$ 9 1.5 $\pm$ 0.3 1.4 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.7 $\pm$ 0.2 2.1 $\pm$ 0.1
4 $\rightarrow$ 10 2.2 $\pm$ 0.3 2.1 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.7 $\pm$ 0.3 2.6 $\pm$ 0.3
5 $\rightarrow$ 1 2.0 $\pm$ 0.3 2.1 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.6 $\pm$ 0.2 2.1 $\pm$ 0.3
5 $\rightarrow$ 2 1.7 $\pm$ 0.3 1.6 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
5 $\rightarrow$ 3 1.7 $\pm$ 0.3 1.7 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
5 $\rightarrow$ 4 1.4 $\pm$ 0.1 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
5 $\rightarrow$ 6 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.3 3.0 $\pm$ 0.9 1.7 $\pm$ 0.3 1.6 $\pm$ 0.4
5 $\rightarrow$ 7 1.8 $\pm$ 0.3 1.8 $\pm$ 0.3 2.3 $\pm$ 0.3 5.9 $\pm$ 2.7 2.2 $\pm$ 0.2 2.4 $\pm$ 0.6
5 $\rightarrow$ 8 1.5 $\pm$ 0.2 1.7 $\pm$ 0.2 2.0 $\pm$ 0.2 1.5 $\pm$ 0.2 2.1 $\pm$ 0.2 2.8 $\pm$ 0.7
5 $\rightarrow$ 9 2.1 $\pm$ 0.3 2.0 $\pm$ 0.3 2.0 $\pm$ 0.2 2.6 $\pm$ 0.1 1.8 $\pm$ 0.3 2.1 $\pm$ 0.1
5 $\rightarrow$ 10 2.6 $\pm$ 0.3 2.5 $\pm$ 0.3 3.1 $\pm$ 0.6 5.3 $\pm$ 0.6 2.8 $\pm$ 0.3 2.6 $\pm$ 0.3
6 $\rightarrow$ 1 2.0 $\pm$ 0.3 2.0 $\pm$ 0.3 2.5 $\pm$ 0.3 5.1 $\pm$ 4.6 2.7 $\pm$ 0.3 2.1 $\pm$ 0.3
6 $\rightarrow$ 2 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3 2.1 $\pm$ 0.2 1.5 $\pm$ 0.3 2.1 $\pm$ 0.2 1.4 $\pm$ 0.4
6 $\rightarrow$ 3 1.5 $\pm$ 0.3 1.5 $\pm$ 0.3 2.3 $\pm$ 0.3 1.2 $\pm$ 0.2 2.2 $\pm$ 0.3 1.6 $\pm$ 0.3
6 $\rightarrow$ 4 1.2 $\pm$ 0.2 1.2 $\pm$ 0.2 2.1 $\pm$ 0.2 1.0 $\pm$ 0.1 2.3 $\pm$ 0.2 1.1 $\pm$ 0.2
6 $\rightarrow$ 5 1.6 $\pm$ 0.2 1.5 $\pm$ 0.2 1.7 $\pm$ 0.1 3.5 $\pm$ 1.5 1.7 $\pm$ 0.2 5.5 $\pm$ 1.0